OpenStudy (anonymous):

Solve the compound inequality 6>-4x+5 or 9<-5x+2 -

OpenStudy (anonymous):

\[6>-4x+5\] subtract 5 \[1>-4x\] divide by -4 don't forget to change the 'sense' of the inequality \[-\frac{1}{4}<x\] OR \[9<-5x+2\] \[7<-5x\] \[-\frac{7}{5}>x\] answer: \[x<-\frac{7}{5}\] or \[x>-\frac{1}{4}\]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

welcome